OpenStudy (anonymous):

The perimeter of a face of a cube is 16 mm. What is its volume? 64 mm3 128 mm3 512 mm3 4096 mm3

OpenStudy (happykiddo):

the first one

OpenStudy (anonymous):

do you need more help?

OpenStudy (phi):

The perimeter of the face is the same as the perimeter of a square: you add up the lengths of each side. There are four equal sides that add up to 16. 4S=16 (as an equation) or if you like S+S+S+S=16 either way, S=4 the volume of a cube is \(S^3\) or S*S*S. So volume is 4*4*4= 16*4= 64 \( mm^3 \)